OpenStudy (anonymous):

y''+4y=t^(2)+3e^t find the general solution of the nonhomogenous differential equation

OpenStudy (amistre64):

find the homogenous part then add to it a guessed at nonhomogenous (particular) part

OpenStudy (amistre64):

what is our characteristic equation to root out?

OpenStudy (amistre64):

its gonna be complex im sure ....

OpenStudy (amistre64):

r^2 + 4 = 0 r = 0 +- 2i right?

OpenStudy (amistre64):

\[y_h=c_1cos(2x)+c_2sin(2x)\] then we have to form a nonhomogenous from the form given above

OpenStudy (amistre64):

and x=t ...

OpenStudy (amistre64):

given: t^(2)+3e^t \[y_p=(Ax^2+B)+(Ce^t)\] since none of these are a copy of what we found before, there doesnt need to be any adjustments

OpenStudy (amistre64):

\[y_p=Ax^2+B+Ce^t\] \[(y_p)'=2Ax+Ce^t\] \[(y_p)''=2A+Ce^t\] \[(y_p)''+4(y_p)=2A+Ce^t+4(Ax^2+B+Ce^t)=(1)t^2+(3)e^t\] simplify and compare coeefs

OpenStudy (turingtest):

\[y_p=Ax^2+Bx+C+De^t\]actually

OpenStudy (amistre64):

helps if i keep my ts in place instead of introducing xs tho ... \[2A+Ce^t+4Ax^2+4B+4Ce^t=(1)t^2+(3)e^t\] \[(4A)t^2=1t^2\] \[5Ce^t=(3)e^t\] \[(2A+4B)=0 \] we got this in order?

OpenStudy (amistre64):

A = 1/4 C = 3/5 2(1/4) + 4B = 0 4B = -1/2 B = -1/8

OpenStudy (amistre64):

... so much typing and ugh ... yeah, x^2+x+k

OpenStudy (amistre64):

i still have to do these on paper lol

OpenStudy (turingtest):

you got the right answer though haha we both have x's and t's all mixed together

OpenStudy (amistre64):

once its done correctly :) \[Y=y_h+y_p\]

OpenStudy (anonymous):

thanks :)